OpenStudy (anonymous):

Write the expression as a single logarithm. Express powers as factors. **open question to see the rest of the problem**

OpenStudy (anonymous):

\[(a.) \log_{6} \sqrt{x}-\log_{6}x^3 \] \[(b.) 4\log_{a}(6x^8)-\frac{ 1 }{ 3 }\log_{a}(8x+15) \]

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

for the first one, you'll use the rule \[\Large \log_{b}(x)-\log_{b}(y)=\log_{b}\left(\frac{x}{y}\right)\]

jimthompson5910 (jim_thompson5910):

so what do you get when you use that for part a?

OpenStudy (anonymous):

\[\frac{ \sqrt{x} }{ x^3}\]

OpenStudy (anonymous):

no wait it'd be \[\log_{6} \frac{ \sqrt{x} }{ x^3 }\]

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

for part b, you have to use this rule first \[\Large y\log_{b}(x)=\log_{b}(x^{y})\]

jimthompson5910 (jim_thompson5910):

and then you can use the rule you used on part a)

OpenStudy (anonymous):

I put in \[\log_{6}\frac{ \sqrt{x} }{ x^3 } \] and it says it's wrong...

jimthompson5910 (jim_thompson5910):

ok try to simplify \[\Large \frac{\sqrt{x}}{x^3}\]

jimthompson5910 (jim_thompson5910):

use the idea \[\Large \sqrt{x} = x^{1/2}\] and subtract exponents

OpenStudy (anonymous):

what do you mean "subtract the exponents"?

jimthompson5910 (jim_thompson5910):

you'll have \[\Large \frac{\sqrt{x}}{x^3} = \frac{x^{1/2}}{x^3}\]

jimthompson5910 (jim_thompson5910):

subtract the exponents 1/2 and 3

OpenStudy (anonymous):

ohhh so that's be -5/2?

OpenStudy (anonymous):

@jim_thompson5910